OpenStudy (anonymous):

What is 2e^4x minus 2e^-4x?

myininaya (myininaya):

how to factor it? if you want to factor it is e(e^x-e^(-2x))(e^x+e^(-2x))

myininaya (myininaya):

oops that first e should be a 2

OpenStudy (anonymous):

no how to subtract it.

myininaya (myininaya):

u cant

OpenStudy (anonymous):

Then what do you do?

myininaya (myininaya):

what do u mean

myininaya (myininaya):

are you trying to show it is equal to something?

OpenStudy (anonymous):

I got it when i was try to solve this matrix inverse problem Find the inverse of the matrix if it exist. |e^x -e^2x| |e^2x e^3x| It may be wrong though...

myininaya (myininaya):

i get 1/(2e^(4x))*|e^3x e^2x| |-e^2x e^x|

OpenStudy (anonymous):

I have that too. but how do i multiply to get the answer?

myininaya (myininaya):

so you want to multiply the scalar inside the matrix

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok i get |(1/(2e^x) 1/(2e^x)| |-1/(2e^(2x)) 1/(2e^3x)|

myininaya (myininaya):

oops error! the first row first column should be 1/(2e^(2x))

myininaya (myininaya):

first row and 2nd column*

myininaya (myininaya):

you just multiply each entry by the scalar

OpenStudy (anonymous):

So those are the final answers?

myininaya (myininaya):

yes with correction i mentioned

OpenStudy (anonymous):

Thank you so much!!!!

myininaya (myininaya):

do you understand how to multiply each entry by the scalar?

myininaya (myininaya):

you said that was your problem?

OpenStudy (anonymous):

yes because of the exponentials. If it was regular numbers I could do it.

myininaya (myininaya):

since the bases are the same law of exponents tell us to subtract the bottom exponent from the top exponent if we get a negative, usually people write it as a positive by flipping it for example (e^(2x)/e^(6x))=e^(2x-6x)=e^(-4x)

myininaya (myininaya):

=1/(e^(4x))

myininaya (myininaya):

another example: (e^(6x)/e^(2x))=e^(6x-2x)=e^(4x)

OpenStudy (anonymous):

ok i see. thank you! i really appriciate you taking the time to show me.

myininaya (myininaya):

np if you want i guess you can be my fan :)